OpenStudy (anonymous):

One solution of the equation 2 x^2 - x - 10 = 0 is 5/2 . What is the other solution?

OpenStudy (anonymous):

the product of the solutions is the constant term...[HINT]

OpenStudy (anonymous):

wait what?

OpenStudy (anonymous):

do you know what is the constant term in your equation is?

OpenStudy (anonymous):

the constant term is the "c" in \(\large ax^2+bx+c=0 \)

OpenStudy (anonymous):

ohhhh thanks

OpenStudy (anonymous):

yw...:)

OpenStudy (anonymous):

so do you have the other solution?

OpenStudy (anonymous):

no

OpenStudy (radar):

Try a -2 for the x value and see what it is equal to.

OpenStudy (radar):

\[2(-2^{2})- (-2) -10=0\] 8 + 2 -10 =0 0=0 What does that tell you.

OpenStudy (radar):

Earlier it was mentioned that the solution was a product of the constant C 2(5/2) times what equals -10 5x=-10 x=-10/5=-2